# Particles before symmetry

### 05 Mar 2026 v5

@@ 5.4 Fundamental bundles and the Standard Model
From the standpoint of the PFB-POV, this is a genuine ambiguity: one can take either the product or the quotient as the structure group, and only global topological data or further physical input can decide between them. From the standpoint of the VB-POV, by contrast, there is nothing to decide. One begins with three fundamental vector bundles—fibres $\mathbb{C}^{3}$, $\mathbb{C}^{2}$, and $\mathbb{C}$, with Hermitian inner products and, for the first two, orientations. Their automorphism groups are $SU(3)$, $SU(2)$, and $U(1)$, and the reconstructed gauge group is their product. The $\mathbb{Z}_{6}$ kernel is a downstream consequence of the specific tensor products nature happens to employ—a curiosity about which multiplets appear in the Standard Model, not a structural problem for the formalism. The analogy is with $SO(3)$ acting on $\mathbb{R}^{3}$: certain tensors built from $\mathbb{R}^{3}$ may be insensitive to certain rotations, but this does not make $SO(3)$ ‘too big’ for the vector space.
@@
Recall the slack world of Section 5.1.a. That is not our world. And barring the standard treatment of the vacuum Yang-Mills case, nor is such a world usually included in the menu of physical possibilities. We do not often seriously consider a mismatch between the symmetries of the gauge sector and those of the matter sector. I claim that there is a tacit assumption built into actual practice. Namely, that each factor of $G$ coincides with the automorphism group of a fundamental bundle:
@@
This is precisely what the VB-POV makes explicit.
@@ 6.1 Summary
The geometry-first formulation does three things. It posits a different ontology. It offers independent, symmetry-free explanations of familiar mechanisms. And it eliminates the slack between symmetry and geometry that the principal-bundle formulation tolerates. That tighter fit may help explain why certain group–geometry correspondences are realised in our world and others are not.
@@
Consider how geometry does the explanatory work of the Higgs mechanism without mentioning symmetry. Since we represent non-isomorphic sections directly from the start, Goldstone modes never appear and never require elimination (see footnote 18). What the symmetry-first formulation calls ‘mass acquisition’ is simply this: the kinetic term of any section depends on the bundle’s affine structure, and the shape operator along the Higgs direction—the extrinsic curvature—is the part that acquires mass. [^35] In the linearised regime, no Killing form is needed: the inner product on vector bundles suffices to give dimensions to perturbations of flat covariant derivatives. But the covariant derivative along a single section does not exhaust the affine degrees of freedom (when $\text{dim}(E_{x})\geq 2$). The degrees absent from the extrinsic curvature remain massless—corresponding, in symmetry-first language, to the unbroken gauge group and its massless photons. Strictly speaking, ‘mass acquisition’ is a misnomer from the geometry-first standpoint; what could it mean for a covariant derivative to acquire mass? But the phrase acquires clear meaning in the linearised regime. [^36] Some would hesitate to say gravitons acquire mass merely because a spacetime admits a kinetic term for a constant-norm vector field; yet that they do is the consensus for such theories (\@BBOPcitep\@BAP\@BBN(Jacobson, 2008)\@BBCP).
The Yukawa mechanism, too, gains clarity. The Yukawa form becomes geometrically natural—a fibrewise contraction built from structures already present on the fundamental bundles—rather than a representation-theoretic construction that postulates maps between associated bundles and absorbs ambiguities into coupling constants. And the formulation sharpens physical questions: the quark Yukawa term depends essentially on the orientation of $\mathbb{C}^{2}$, which explains why $SU(2)$ rather than $U(2)$ appears naturally. For $\mathbb{C}^{3}$, by contrast, I don’t know the analogous mechanism—why does the Standard Model employ $SU(3)$ rather than U(3)U(3)? [^37] There is a canonical isomorphism $U(3)\simeq(SU(3)\times U(1))/\mathbb{Z}^{3}$, but the $U(1)$ representations in the Standard Model do not realise it. Benjamin Muntz (p.c.) suggests looking at triality constraints on baryon coupling: colourless three-quark states would not be invariant under $U(3)$.
@@ 6.2 Two defences
The geometry-first picture has comparative advantages. The symmetry-free explanations just surveyed—of the Higgs mechanism, the Yukawa couplings, and charge quantisation—share a common source: the geometry-first picture eliminates the slack between symmetry and geometry that the principal-bundle picture tolerates. It makes explicit an assumption—that gauge groups coincide with automorphism groups of fundamental bundles—that the principal-bundle picture leaves tacit, and thereby excludes many symmetry-first models—most of which are never considered in practice anyway. [^38] As discussed in Section 5, there is a widespread tacit assumption that gauge groups simply are automorphism groups of fundamental vector spaces. But this assumption is both unwarranted and inconsistently applied.
This is a defence of the geometry-first picture in absolute terms: I judge it to provide superior explanations—a cleaner ontology—for the class of theories to which it applies. The principal gauge field is not altogether denied reality; in the geometry-first picture, the same physical content is carried by the covariant derivative on a specific vector bundle. But I acknowledge that my preference is staked on ontological commitments that not everyone shares.
Of course, the geometry-first picture also has comparative disadvantages. It cannot accommodate gauge theories built from exceptional Lie groups such as $E_{8}$ without losing its explanatory force (Section 5.2). Moreover, the principal-bundle picture’s greater flexibility is a genuine virtue when one wants a single framework covering as wide a class of gauge theories as possible, not just the narrower class that includes the Standard Model.
So let me offer a second defence that does not require my aforementioned commitments. What we can say unconditionally is that, for the Standard Model, the predictions stay the same. Why bother, then, with a reformulation that changes no predictions? As Feynman observed, and as \@BBOPcitet\@BAP\@BBNHunt (2025)\@BBCP has argued in detail, equivalent formulations are not always equally fertile. A genuinely different explanation of the same phenomena is already an epistemic gain—it widens the base from which we move into the unknown. The comparative advantages noted above are sufficient to warrant the geometry-first picture’s consideration as an alternative, even if it does not supersede the principal-bundle picture for all purposes.
@@ Appendix A Principal and associated fibre bundles
Given an element $\xi$ of the Lie-algebra $\mathfrak{g}$, and the action of $G$ on $P$, we use the exponential to find an action of $\mathfrak{g}$ on $P$. This defines an embedding of the Lie algebra into the tangent space at each point, given by the hash operator: $\sharp_{p}:\mathfrak{g}\rightarrow T_{p}P$. The image of this embedding we call the vertical space $V_{p}$ at a point $p\in P$: it is tangent to the orbits of the group, and is linearly spanned by vectors of the form
@@
Definition 2 (An principal connection-form)
$\varpi$ is defined as a Lie-algebra valued one form on $P$, satisfying the following properties:
@@
Now, in possession of an principal connection, we can induce a notion of covariant derivative on associated vector bundles:
@@
The existence and uniqueness of this map is guaranteed for $\gamma\subset U$ some open subset of $M$, and it follows from properties of solutions of ordinary differential equations (cf. \@BBOPcitep\@BAP\@BBN(Kobayashi and Nomizu, 1963, Ch. II.2)\@BBCP).

### 23 Mar 2026 v6

@@ 5.4 Fundamental bundles and the Standard Model
From the standpoint of the PFB-POV, this is a genuine ambiguity: one can take either the product or the quotient as the structure group, and only global topological data or further physical input can decide between them. (Indeed, the quotient $G/\mathbb{Z}_{6}$ acts faithfully on all matter yet cannot be produced by any geometry-first starting point, since the $\mathbb{Z}_{6}$ sits diagonally across the three factors; this is the kernel-entanglement obstruction of \@BBOPcitet\@BAP\@BBNGomes (2026)\@BBCP.) From the standpoint of the VB-POV, by contrast, there is nothing to decide. One begins with three fundamental vector bundles—fibres $\mathbb{C}^{3}$, $\mathbb{C}^{2}$, and $\mathbb{C}$, with Hermitian inner products and, for the first two, orientations. Their automorphism groups are $SU(3)$, $SU(2)$, and $U(1)$, and the reconstructed gauge group is their product. The $\mathbb{Z}_{6}$ kernel is a downstream consequence of the specific tensor products nature happens to employ—a curiosity about which multiplets appear in the Standard Model, not a structural problem for the formalism. The analogy is with $SO(3)$ acting on $\mathbb{R}^{3}$: certain tensors built from $\mathbb{R}^{3}$ may be insensitive to certain rotations, but this does not make $SO(3)$ ‘too big’ for the vector space.
@@
Recall the slack world of Section 5.1.b. That is not our world. And barring the standard treatment of the vacuum Yang-Mills case, nor is such a world usually included in the menu of physical possibilities. We do not often seriously consider a mismatch between the symmetries of the gauge sector and those of the matter sector. I claim that there is a tacit assumption built into actual practice. Namely, that each factor of $G$ coincides with the automorphism group of a fundamental bundle:
@@
This is precisely what the VB-POV makes explicit. Yet no standard textbook presents the Standard Model matter content as arising tensorially from fundamental vector spaces. Every treatment I know—\@BBOPcitet\@BAP\@BBNPeskin & Schroeder (1995)\@BBCP, \@BBOPcitet\@BAP\@BBNWeinberg (1996)\@BBCP, \@BBOPcitet\@BAP\@BBNBleecker (1981)\@BBCP, \@BBOPcitet\@BAP\@BBNHamilton (2017)\@BBCP, \@BBOPcitet\@BAP\@BBNNakahara (2003)\@BBCP, \@BBOPcitet\@BAP\@BBNTong (2025)\@BBCP—follows the same pattern: postulate the gauge group, then list the representations as empirical input. Knowing that representations of classical groups arise from tensor products of fundamental representations is not the same as building that fact into the formalism.
@@ 6.1 Summary
The geometry-first formulation does three things. It posits a different ontology. It offers independent, symmetry-free explanations of familiar mechanisms. And it structures the relationship between symmetry and geometry more tightly than the principal-bundle formulation requires—demanding that each gauge factor arise as the automorphism group of a fundamental bundle. That tighter fit may help explain why certain group–geometry correspondences are realised in our world and others are not.
@@
Consider how geometry does the explanatory work of the Higgs mechanism without mentioning symmetry. Since we represent non-isomorphic sections directly from the start, Goldstone modes never appear and never require elimination (see footnote 15). What the symmetry-first formulation calls ‘mass acquisition’ is simply this: the kinetic term of any section depends on the bundle’s affine structure, and the shape operator along the Higgs direction—the extrinsic curvature—is the part that acquires mass. [^32] In the linearised regime, no Killing form is needed: the inner product on vector bundles suffices to give dimensions to perturbations of flat covariant derivatives. But the covariant derivative along a single section does not exhaust the affine degrees of freedom (when $\text{dim}(E_{x})\geq 2$). The degrees absent from the extrinsic curvature remain massless—corresponding, in symmetry-first language, to the unbroken gauge group and its massless photons. Strictly speaking, ‘mass acquisition’ is a misnomer from the geometry-first standpoint; what could it mean for a covariant derivative to acquire mass? But the phrase acquires clear meaning in the linearised regime. [^33] Some would hesitate to say gravitons acquire mass merely because a spacetime admits a kinetic term for a constant-norm vector field; yet that they do is the consensus for such theories (\@BBOPcitep\@BAP\@BBN(Jacobson, 2008)\@BBCP).
The Yukawa mechanism, too, gains clarity. The Yukawa form becomes geometrically natural—a fibrewise contraction built from structures already present on the fundamental bundles—rather than a representation-theoretic construction that postulates maps between associated bundles and absorbs ambiguities into coupling constants. And the formulation sharpens physical questions: the quark Yukawa term depends essentially on the orientation of $\mathbb{C}^{2}$, which explains why $SU(2)$ rather than $U(2)$ appears naturally. For $\mathbb{C}^{3}$, by contrast, I don’t know the analogous mechanism—why does the Standard Model employ $SU(3)$ rather than U(3)U(3)? [^34] There is a canonical isomorphism $U(3)\simeq(SU(3)\times U(1))/\mathbb{Z}_{3}$, but the $U(1)$ representations in the Standard Model do not realise it. Benjamin Muntz (p.c.) suggests looking at triality constraints on baryon coupling: colourless three-quark states would not be invariant under $U(3)$.
@@ 6.2 Two defences
I take the geometry-first picture to provide superior explanations—a cleaner ontology—for the class of theories to which it applies. The principal gauge field is not denied reality; the same physical content is carried by the covariant derivative on a vector bundle. But this preference rests on ontological commitments that not everyone shares.
The geometry-first picture also has clear disadvantages. It cannot accommodate gauge theories built from exceptional Lie groups such as $E_{8}$ without losing its explanatory force (Section 5.2), and the principal-bundle picture’s greater flexibility is a genuine virtue when one wants a single framework covering as wide a class of gauge theories as possible.
So let me offer a second defence that does not require those commitments. For the Standard Model, the predictions stay the same. Why bother, then, with a reformulation that changes no predictions? Feynman made the point in his Nobel Lecture:
The subsequent career of Feynman’s path integrals—equivalent to Schwinger’s operator methods, yet far more fertile—bears this out. So does Minkowski’s 1908 recasting of special relativity, without which general relativity could hardly have been conceived. [^35] As \@BBOPcitet\@BAP\@BBNHunt (2025)\@BBCP argues, equivalent formulations are not always equally fertile. Under Hunt’s lights, a reformulation is non-trivial when it suggests a different ontology, offers new explanations, and displays an epistemic difference in how it solves problems: “Successful reformulations clarify what we need to know to solve problems, improving our understanding of the world” (ibid. p. 5). A genuinely different explanation of the same phenomena widens the base from which we move into the unknown.
Nor should the two formalisms be called ‘equivalent.’ As shown in a companion paper \@BBOPcitep\@BAP\@BBN(Gomes, 2026)\@BBCP, the natural functor $\mathcal{F}$ from the geometry-first category to the category of principal bundles with matter is faithful but neither essentially surjective nor full—even when the codomain is restricted to theories in which the gauge group acts faithfully on the total matter content. Distinct geometry-first data always yield distinct symmetry-first outputs; but legitimate PFB-POV theories exist—including the Standard Model with its $\mathbb{Z}_{6}$-quotiented gauge group—that no geometry-first starting point can produce; and the PFB-POV admits symmetries (outer automorphisms) that the VB-POV structurally cannot generate. Furthermore, the principal bundle together with its matter bundles does not record which bundles are fundamental, nor how the matter representations descend tensorially from them. The two formalisms also differ in the space of possible theories they cover: the VB-POV excludes non-integer charges and gauge groups from exceptional families without low-dimensional faithful representations (Section 5.2). The analogy is with Barbour–Bertotti relational mechanics and Newtonian particle mechanics: the two agree on the zero-angular-momentum sector, yet differ in the models they admit. [^36] See \@BBOPcitet\@BAP\@BBNBarbour & Bertotti (1982)\@BBCP; for philosophical discussion, \@BBOPcitet\@BAP\@BBNPooley & Brown (2002)\@BBCP. Agreement on predictions within a domain should not be mistaken for equivalence of theoretical frameworks.
@@ Appendix A Principal and associated fibre bundles
Given an element $\xi$ of the Lie algebra $\mathfrak{g}$, and the action of $G$ on $P$, we use the exponential to find an action of $\mathfrak{g}$ on $P$. This defines an embedding of the Lie algebra into the tangent space at each point, given by the hash operator: $\sharp_{p}:\mathfrak{g}\rightarrow T_{p}P$. The image of this embedding we call the vertical space $V_{p}$ at a point $p\in P$: it is tangent to the orbits of the group, and is linearly spanned by vectors of the form
@@
Definition 2 (A principal connection-form)
$\varpi$ is defined as a Lie algebra valued one form on $P$, satisfying the following properties:
@@
Now, in possession of a principal connection, we can induce a notion of covariant derivative on associated vector bundles:
@@
The existence and uniqueness of this map is guaranteed for $\gamma\subset U$ some open subset of $M$, and it follows from properties of solutions of ordinary differential equations (cf. \@BBOPcitep\@BAP\@BBN(Kobayashi & Nomizu, 1963, Ch. II.2)\@BBCP).